OpenStudy (anonymous):

(-15x^2 over 12b2) (ab over x^2) these are meant to be fractions

OpenStudy (anonymous):

Welcome to Openstudy! I can help you show your question in terms of fractions. You jus tuse the equation down here in the comments. Anyways this question will turn inot \[\Large \frac{-15x^2}{12b^2} \cdot \frac{ab}{x^2}\] Is this it?